OpenStudy (anonymous):

y = 2x − x2, x = 2, Δx = −0.5. Δy= ? dy= ?

OpenStudy (aum):

\[ y = 2x - x^2 \\ f(x) = 2x - x^2 \\ f(2) = 2*2 - 2^2 = 4 - 4 = 0 \\ f(x+\Delta x) = f(2 - 0.5) = f(1.5) = 2*1.5 - 1.5^2 = 3.0 - 2.25 = 0.75 \\ \Delta y = f(x+\Delta x) - f(x) = 0.75 - 0.0 = 0.75 \\ \]\[ y = 2x - x^2 \\ \frac{dy}{dx} = (2 - 2x) \\ \text{At x = 2: } \\ \frac{dy}{dx} = (2 - 2*2) = -2 \\ dy = -2dx \\ dx \approx \Delta x = -0.5 \\ dy = -2 * (-0.5) = 1 \]